# Unveiling Molecular Moieties through Hierarchical Graph Explainability

### v2

@@ 3.3 Explainability results
As already detailed in Section 2.3, the main contribution of our work is the hierarchical Grad-CAM graph explainability procedure to devise relevant moieties at different scales in the molecule. Two visual examples of the results provided by our hierarchical graph explainer compared to the docking analysis carried out by an expert in Computational Chemistry are shown in Figure 4 and Figure 5. In such figures, two well known drugs acting as CDK1 inhibitors are reported: the Dinaciclib (PubChem ID: 46926350) and the Meridianin e (PubChem ID: 9995236). The color codes are the one already detailed in Section 2.3.
The explainability procedure was carried out on 19 CDK1 inhibitors, and the results have been compared to either the docking results for the same inhibitors or experimental crystal structures of co-crystallised ligands, when available. The complete visual results for all the drugs are reported in the Supplementary Materials. Our methodology is consistent with what was observed using experimental structures in 17 of 19 known drugs, and they are reported in Table 5 which lists the generic name of the drug, the PubChem ID, a binary assessment made by the expert on the goodness of prediction of the molecular features involved in the host-guest interaction between drug and target that we called Guest Feature Prediction (GFP) and the detailed accordance of the moieties coloring to the docking outcomes. Here, the expert verified if each piece of information gathered from the docking had been colored properly by our explainer that is with a label in accordance with the docking procedure. We reported such information in terms of single statements that have been red colored in case the explainer missed the correct label.
@@ 5 Conclusion
The use of AI in chemoinformatics allowed the rise of ligand-based methods that were outclassed by structure-based methods in the last years. Structure-based methods offer the possibility to look deeper in the binding site of a protein allowing scientists to predict how ligand could bind to a specific target. Anyway, It is well known that the use of structure-based methods in drug design could be strongly affected by the protein and ligand conformation in retrieving important protein-ligand interactions. Moreover, high quality protein-ligand complexes are not always available in public databases. In our work, a novel mixed ligand/structure-based VS approach has been presented in this work, which exploits structural information using a novel eXplainable AI (XAI) procedure on a Deep Neural Network trained for a ligand-based VS task. The presented explainer leverages the algorithmic characteristics of the well-known Grad-CAM explainer for Graph Convolutional Neural Networks based on message passing. Moving from the SOTA results obtained in ligand-based VS with a multi-target classifier which uses the EMBER molecular descriptor, we built a novel classifier relying on GCNNs applied to molecular graphs. Such a multi-target classifier outperformed our previous result, and formed the trained backbone for our explainability analysis. The main idea is using a hierarchy of Grad-CAM explainers placed at different layers of the network where the message passing takes into account a wider and wider neighborhood of each atom. The explanations at each level are then summed together to provide a global map where the different molecular moieties involved in the bioactivity prediction are highlighted and labelled with their role using the well known RDKit package. In our implementation, we applied the explainer to a single target version of the classifier to avoid the inherent classification bias arising in multi-class architectures. We built a CDK1-only classifier that was tuned properly, and the explainability results on 19 well known drugs with CDK1 inhibitory action were tested by a Computational Chemistry expert. The results are very good: 17 of 19 explanations are in accordance with the docking analysis carried out by the expert.
@@
Differently from pure structure-based methods, the results obtained from our method enlighten not only the interacting partners but also the molecular moieties involved, for example, in the hydrophobic interactions (useful for entropy driven binding) and other chemical parts responsible for the recognition of the target. The main strength of our method is the use of combined explainability layers that taken together are capable to account for both the whole molecular structure and functional groups connections. As a result of this architecture, it is possible to correctly assign a target to each molecule and retrieve the main molecule moieties involved in the target assignation/binding. Such a result is not retrievable in other computational techniques used for Virtual Screening, letting our method to unveil hidden molecular moieties useful for binding a specific target. Our methods The explainability approach herein presented, will be applied to other targets, and the molecules prioritised by the multi-target classifier will be tested on a panel of kinases in order to experimentally validate the approach.
Supplementary information
@@
Abbrevations
  • CADD. Computer-Aided Drug Design
  • VS. Virtual Screening
  • SAR. Structure Activity Relationship
  • XAI. Explainable AI
  • EMBER. EMBedding multiplE molecular fingeRprints
  • IC50. Half Maximal Inhibitory Concentration
  • Ki. Inhibition constant
  • Kd. Dissociation constant
  • MLP. Multi Layer Perceptron
  • PDB. Protein Data Bank
  • ReLU. Rectified Linear Unit
  • wBCE. Weighted Binary Cross-Entropy
  • Grad-CAM. Gradient-weighted Class Activaction Mapping
  • TP. True Positive
  • P. Positive
  • TN. True Negative
  • N. Negative
  • TP/P. True Positive vs Positive ratio
  • EF. Enrichment Factor

### v3

@@ 3.3 Explainability results
As already detailed in Section 2.3, the main contribution of our work is the hierarchical Grad-CAM graph explainability procedure to devise relevant moieties at different scales in the molecule. Two visual examples of the results provided by our hierarchical graph explainer compared to experimental x-ray structures and the molecular docking pose are shown in Figure 4 and Figure 5. In such figures, two well known drugs acting as CDK1 inhibitors are reported: the Dinaciclib (PubChem ID: 46926350) and the Meridianin e (PubChem ID: 9995236). The color codes are the one already detailed in Section 2.3.
The explainability procedure was carried out on 19 CDK1 inhibitors, and the results have been compared to either the docking results for the same inhibitors or experimental crystal structures of co-crystallised ligands, when available. The complete visual results for all the drugs are reported in the Supplementary Materials. Our methodology is consistent with what was observed using experimental structures in 17 of 19 known drugs, and they are reported in Table 5 which lists the generic name of the drug, the PubChem ID, a binary assessment made by comparison with known data on the goodness of prediction of the molecular features involved in the host-guest interaction between drug and target that we called Guest Feature Prediction (GFP) and the detailed accordance of the moieties coloring to the docking outcomes. Here, the comparison of results by our explainer expressed by labels were evaluated based on accordance with the experimental and docking procedure. We reported such information in terms of single statements that have been red colored in case the explainer missed the correct label.
@@ 5 Conclusion
The use of AI in chemoinformatics allowed the rise of ligand-based methods that were outclassed by structure-based methods in the last years. Structure-based methods offer the possibility to look deeper in the binding site of a protein allowing scientists to predict how ligand could bind to a specific target. Anyway, it is well known that the use of structure-based methods in drug design could be strongly affected by the protein and ligand conformation in retrieving important protein-ligand interactions. Moreover, high quality protein-ligand complexes are not always available in public databases. In our work, a novel mixed ligand/structure-based approach has been presented, which exploits structural information using a novel eXplainable AI (XAI) procedure on a Deep Neural Network trained for a ligand-based VS task. The presented explainer leverages the algorithmic characteristics of the well-known Grad-CAM explainer for Graph Convolutional Neural Networks based on message passing. Moving from the SOTA results obtained in ligand-based VS with a multi-target classifier which uses the EMBER molecular descriptor, we built a novel classifier relying on GCNNs applied to molecular graphs. Such a multi-target classifier outperformed our previous result, and formed the trained backbone for our explainability analysis. The main idea is using a hierarchy of Grad-CAM explainers placed at different layers of the network where the message passing takes into account a wider and wider neighborhood of each atom. The explanations at each level are then summed together to provide a global map where the different molecular moieties involved in the bioactivity prediction are highlighted and then labelled using the well known RDKit package. In our implementation, we applied the explainer to a single target version of the classifier to avoid the inherent classification bias arising in multi-class architectures. We built a CDK1-only classifier that was tuned properly, and the explainability results on 19 well known drugs with CDK1 inhibitory action were compared to literature, experimental and docking data. The results are very good: 17 of 19 explanations are in accordance with the docking analysis.
@@
Differently from pure structure-based methods, the results obtained from our method enlighten not only the interacting partners but also the molecular moieties involved, for example, in the hydrophobic interactions (useful for entropy driven binding) and other chemical parts responsible for the recognition of the target. The main strength of our method is the use of combined explainability layers that taken together are capable to account for both the whole molecular structure and functional groups connections. As a result of this architecture, it is possible to correctly assign a target to each molecule and retrieve the main molecule moieties involved in the target assignation/binding. Such a result is not retrievable in other computational techniques used for Virtual Screening, letting our method to unveil hidden molecular moieties useful for binding a specific target. The explainability approach herein presented, will be applied to other targets, and the molecules prioritised by the multi-target classifier will be tested on a panel of kinases in order to experimentally validate the approach.
6 Data and Code Availability
7 Supplementary information
@@
8 Acknowledgments
This work has ben partially supported by project:
1.“SAMOTHRACE” (Sicilian MicronanoTech Research And Innovation Center), cup project B73C22000810001, project code ECS_00000022.